Moravia Company processes and packages cream cheese. The following data have been compiled for the month of April. Conversion activity occurs uniformly throughout the production process.

Work in process, April 1—13,000 units:
Direct material: 100% complete, cost of $ 20,000
Conversion: 25% complete, cost of 3,500
Balance in work in process, April 1 $ 23,500
Units started during April 112,000
Units completed during April and transferred out to finished-goods inventory 85,000
Work in process, April 30:
Direct material: 100% complete
Conversion: 35% complete
Costs incurred during April:
Direct material $ 230,000
Conversion costs:
Direct labor $ 46,200
Applied manufacturing overhead 84,940
Total conversion costs $ 131,140


Prepare schedules to accomplish each of the following process-costing steps for the month of April. Use the weighted-average method of process costing.


Required:

1. Analysis of physical flow of units.

2. Calculation of equivalent units.

3. Computation of unit costs. (Round "Cost per Equivalent Unit" to 2 decimal places.)
Answer #1

Units in Ending WIP = Units in Beginning WIP +Units started - Units completed and transferred
= 13000 + 112000 - 85000 = 40000 units

1.

Physical Flow Schedule
Units to account for
     Beginning WIP 13000
      Units started 112000
Total Units to account for 125000
Units accounted for
    Units completed and transferred out
Beginning WIP 13000
Units started 72000 85000
Ending WIP 40000
Total Units accounted for 125000

2. Equivalent units of Production = Units completed and transferred + Ending WIP x Completion
Material = 85000 + 40000 x 100% = 125000 units
Conversion = 85000 + 40000 x 35% = 99000 units

3. Cost per equivalent unit = (Beginning Cost + Cost added) / Equivalent units
Material = ($20000 + $230000) / 125000 = $2 per unit
Conversion = ($23500 + $131140) / 99000 = $1.56 per unit

Total Cost per unit = $3.56 per unit